3. A firm sells a product for birr 60 per unit. Raw material cost are birr 13 per unit, laborcosts are birr 20 per


unit, and annual fixed costs are birr 340,000.


(a) Determine the profit function.


(b) How many units would have to sold to earn an annual profit of birr 240,000?


(c) Determine the amount of contribution to the fixed cost and profit, if the firm sells


10,000units.

1
Expert's answer
2022-04-06T14:47:32-0400

a.

"Profit=total \space sales -total \space costs"

Let X be units sold

"Px=60x-(13+20)x-340000\\Px=60x-33x-340000\\
Px=27x-340000"

b.

"27x-340000,=240000\\27x=240000+340000\\27x=580000\\x=21481.5units"

c.

"Px=27x-340000\\=27(10000)-340000\\=70000 loss"


"%" annual fixed costs